### v2

# A note on the omega-chaos

Abstract.
For any continuous self-map of a compact metric space, we provide sufficient conditions under which the infinite direct product of the map is $\omega$-chaotic. We also apply the result to obtain some examples of unusual $\omega$-chaotic maps.
Key words and phrases:
continuous maps, direct products, $\omega$-chaos, $\omega^{\ast}$-chaos, proximal
2020 Mathematics Subject Classification:
37B05, 37D45
Chaos is a key concept in the modern theory of dynamical systems and is relevant to unpredictable phenomena that can occur even in deterministic systems. One mathematical definition of chaos is Li–Yorke chaos, which focuses on the complicated long-term behavior of pairs of orbits [8]. A related definition is $\omega$-chaos that highlights the intricate intersections of $\omega$-limit sets [7]. Since its introduction by S. Li in the context of interval maps, $\omega$-chaos and its variants have been the subject of extensive studies, revealing subtle structures in the asymptotic behavior of dynamical systems.
In this paper, we provide some criteria for $\omega$-chaos and examine its relationship with other dynamical properties. Specifically, for any continuous self-map of a compact metric space, we give sufficient conditions under which the infinite direct product of the map is $\omega$-chaotic. We also apply the result to obtain some examples of unusual $\omega$-chaotic maps, e.g., an $\omega$-chaotic map that is proximal and so not $\omega^{\ast}$-chaotic.
We recall the definition of $\omega$-chaos from [7]. Throughout, $X$ denotes a compact metric space endowed with a metric $d\colon X\times X\to[0,\infty)$. Given a continuous map $f\colon X\to X$ and $x\in X$, we denote by $\omega(x,f)$ the $\omega$-limit set of $x$ for $f$, i.e., the set of $y\in X$ such that
$\lim_{j\to\infty}f^{i_{j}}(x)=y$
for some $0\leq i_{1}<i_{2}<\cdots$. Note that $\omega(x,f)$ is a closed subset of $X$ and satisfies $f(\omega(x,f))=\omega(x,f)$.
Definition 1.
Let $f\colon X\to X$ be a continuous map. We denote by $Per(f)$ the set of periodic points for $f$:
$Per(f)=\bigcup_{j\geq 1}\{x\in X\colon f^{j}(x)=x\}.$
A subset $S$ of $X$ is said to be an $\omega$-scrambled set for $f$ if for any $x,y\in S$ with $x\neq y$,
  • $\omega(x,f)\setminus\omega(y,f)$ is an uncountable set,
  • $\omega(x,f)\cap\omega(y,f)\neq\emptyset$,
  • $\omega(x,f)\setminus Per(f)\neq\emptyset$.
We say that $f$ is $\omega$-chaotic if there is an uncountable $\omega$-scrambled set for $f$.
The main result of this paper is the following theorem.
Theorem 1.
Given a continuous map $f\colon X\to X$, let $X^{\mathbb{N}}$ be the product space of infinitely many copies of $X$ and let
$g=f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$
be the map defined by: for any $u=(u_{n})_{n\geq 1},v=(v_{n})_{n\geq 1}\in X^{\mathbb{N}}$, $v=g(u)$ if and only if $v_{n}=f(u_{n})$ for all $n\geq 1$. If there are a closed subset $\Lambda$ of $X$ with $f(\Lambda)\subset\Lambda$, $p\in\Lambda$, and $z\in X$ such that
  • $\omega((p,z),f\times f)\cap\Delta\neq\emptyset$ where $\Delta=\{(x,y)\in X\times X\colon x=y\}$,
  • $\omega(z,f)\setminus\Lambda$ is an uncountable set,
  • $\omega(z,f)\setminus Per(f)\neq\emptyset$,
then $g$ is $\omega$-chaotic.
Before proving Theorem 1, we state and prove two corollaries of Theorem 1.
Corollary 1.
Given a continuous map $f\colon X\to X$ and $x\in X$, let
$Y=\{f^{i}(x)\colon i\geq 0\}\cup\omega(x,f),$
$h=f|_{Y}\colon Y\to Y$, and let $g=h^{\mathbb{N}}\colon Y^{\mathbb{N}}\to Y^{\mathbb{N}}$. If
  • $\omega(x,f)$ is an uncountable set,
  • $\omega(x,f)\cap Per(f)\neq\emptyset$,
  • $\omega(x,f)\setminus Per(f)\neq\emptyset$,
then $g$ is $\omega$-chaotic.
Proof.
Let $p\in\omega(x,f)\cap Per(f)$ and take $j\geq 1$ such that
  • $f^{j}(p)=p$,
  • $f^{i}(p)\neq p$ for all $1\leq i\leq j-1$.
Since
$p\in\omega(x,f)=\bigcup_{i=0}^{j-1}\omega(f^{i}(x),f^{j}),$
we have $p\in\omega(f^{i}(x),f^{j})$ for some $0\leq i\leq j-1$. Let $z=f^{i}(x)$ and
$\Lambda=\{f^{k}(p)\colon 0\leq k\leq j-1\}.$
Since $f^{j}(p)=p$ and $p\in\omega(z,f^{j})$, we have
$(p,p)\in\mathbb{\omega}((p,z),f\times f);$
therefore, $\omega((p,z),f\times f)\cap\Delta\neq\emptyset$. Since $\omega(z,f)=\omega(x,f)$,
  • $\omega(z,f)\setminus\Lambda$ is an uncountable set,
  • $\omega(z,f)\setminus Per(f)\neq\emptyset$,
thus, by Theorem 1, we conclude that $g$ is $\omega$-chaotic. ∎
A continuous map $f\colon X\to X$ is said to be transitive if for any non-empty open subsets $U,V$ of $X$, there is $i\geq 0$ such that $f^{i}(U)\cap V\neq\emptyset$. Note that $f$ is transitive if and only if $X=\omega(x,f)$ for some $x\in X$. A non-empty closed subset $M$ of $X$ with $f(M)\subset M$ is called a minimal set for $f$ if $M=\omega(x,f)$ for all $x\in M$. We say that $f$ is non-minimal if $X$ is not a minimal set for $f$.
Corollary 2.
Let $f\colon X\to X$ be a continuous map and let $g=f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$. If $f$ is transitive and non-minimal, then $g$ is $\omega$-chaotic.
Proof.
Since $f$ is transitive, we have $X=\omega(z,f)$ for some $z\in X$. By the Auslander–Ellis theorem (see, e.g., [1, Theorem 5]), there are a minimal set $M$ for $f$ and $p\in M$ such that
$\liminf_{i\to\infty}d(f^{i}(p),f^{i}(z))=0$
and so
$\omega((p,z),f\times f)\cap\Delta\neq\emptyset.$
Note that $X\setminus M\neq\emptyset$ because $f$ is non-minimal. Since $X(=\omega(z,f))$ is perfect, i.e., has no isolated point, $X\setminus M$ is perfect and locally compact Hausdorff, thus $X\setminus M(=\omega(z,f)\setminus M)$ is uncountable. Since
$\omega(z,f)\setminus Per(f)=X\setminus Per(f)\neq\emptyset,$
letting $\Lambda=M$, we see that the assumptions of Theorem 1 are satisfied; therefore, $g$ is $\omega$-chaotic. ∎
We shall prove Theorem 1. For the proof, we recall a simplified version of the so-called Kuratowski–Mycielski theorem [9]. For a topological space $Z$, a subset $S$ of $Z$ is called a $G_{\delta}$-subset of $Z$ if $S$ is a countable intersection of open subsets of $Z$. By the Baire category theorem, if $Z$ is a complete metric space, then every countable intersection of dense $G_{\delta}$-subsets of $Z$ is dense in $Z$.
Theorem (Kuratowski–Mycielski).
Let $P$ be a perfect complete separable metric space. If $R$ is a dense $G_{\delta}$-subset of $P\times P$, then there is a Cantor set $S$ in $P$ such that
$S\times S\setminus\Delta\subset R,$
where $\Delta=\{(x,y)\in P\times P\colon x=y\}$.
By this theorem, we obtain the following lemma.
Lemma 1.
Consider $\{p,z\}$ (with the discrete topology) and the product space
$P=\{p,z\}^{\mathbb{N}}.$
Then, there is a Cantor set $S$ in $P$ such that for all $s=(s_{n})_{n\geq 1},t=(t_{n})_{n\geq 1}\in S$, $s\neq t$ implies $(s_{m},t_{m})=(z,p)$ for some $m\geq 1$.
Proof.
Note that $P$ is a perfect compact metrizable space. Let $R$ be the set of
$(s,t)=((s_{n})_{n\geq 1},(t_{n})_{n\geq 1})\in P\times P$
such that $(s_{m},t_{m})=(z,p)$ for some $m\geq 1$. We easily see that $R$ is a dense open subset of $P\times P$. By the above theorem, we obtain a Cantor set $S$ in $P$ such that for all $\in S, sts\neq t$ implies $(s,t)\in R$, i.e., $(s_{m},t_{m})=(z,p)$ for some $m\geq 1$; therefore, the lemma has been proved. ∎
By using Lemma 1, we prove Theorem 1.
Proof of Theorem 1.
As $\{p,z\}\subset X$, we have $\{p,z\}^{\mathbb{N}}\subset X^{\mathbb{N}}$. By
$\omega((p,z),f\times f)\cap\Delta\neq\emptyset,$
we obtain $(r,r)\in\omega((p,z),f\times f)$ for some $r\in\Lambda$. It follows that
$(r,r,r,\dots)\in\omega(s,g)$
for each $s\in\{p,z\}^{\mathbb{N}}$; therefore $\omega(s,g)\cap\omega(t,g)\neq\emptyset$ for all $s,t\in\{p,z\}^{\mathbb{N}}$. Let $s=(s_{n})_{n\geq 1},t=(t_{n})_{n\geq 1}\in\{p,z\}^{\mathbb{N}}$ and assume that $(s_{m},t_{m})=(z,p)$ for some $m\geq 1$. Consider the map
$\pi_{m}\colon X^{\mathbb{N}}\to X$
defined by $\pi_{m}(u)=u_{m}$ for all $u=(u_{n})_{n\geq 1}\in X^{\mathbb{N}}$. For any $w\in\omega(z,f)$, we have
$(q,w)\in\omega((p,z),f\times f)$
for some $q\in\Lambda$. We define $v=(v_{n})_{n\geq 1}(\in X^{\mathbb{N}})$ by
$v_{n}=\begin{cases}q&\text{if $s_{n}=p$}\\ w&\text{if $s_{n}=z$}\end{cases}$
for all $n\geq 1$. It follows that $v\in\omega(s,g)$ and $\pi_{m}(v)=w$. As $w\in\omega(z,f)$ is arbitrary, we obtain $\omega(z,f)\subset\pi_{m}(\omega(s,g))$. Since $t_{m}=p$, we have $\pi_{m}(\omega(t,g))\subset\Lambda$. By these conditions, we obtain
$\omega(z,f)\setminus\Lambda\subset\pi_{m}(\omega(s,g)\setminus\omega(t,g)).$
As $\omega(z,f)\setminus\Lambda$ is uncountable, so is $\omega(s,g)\setminus\omega(t,g)$. Note that $\pi_{m}$ satisfies
$\pi_{m}\circ g=f\circ\pi_{m}.$
If $\omega(s,g)\setminus Per(g)=\emptyset$, then by $\omega(z,f)\subset\pi_{m}(\omega(s,g))$, we obtain $\omega(z,f)\setminus Per(f)=\emptyset$, a contradiction. It follows that $\omega(s,g)\setminus Per(g)\neq\emptyset$. By Lemma 1, we can take a Cantor set $S$ in $\{p,z\}^{\mathbb{N}}$ such that for all $s=(s_{n})_{n\geq 1},t=(t_{n})_{n\geq 1}\in S$, $s\neq t$ implies $(s_{m},t_{m})=(z,p)$ for some $m\geq 1$. Since $S$ is an $\omega$-scrambled set for $g$, we conclude that $g$ is $\omega$-chaotic, completing the proof of the theorem. ∎
By using Corollary 1, we will give several examples of $\omega$-chaotic maps with additional dynamical properties. For this purpose, we recall the notion of abstract $\omega$-limit sets and related facts [2].
Given a continuous map $f\colon X\to X$ and $\delta>0$, a finite sequence $(x_{i})_{i=0}^{k}$ of points in $X$, where $k\geq 1$, is called a $\delta$-chain of $f$ if
$\sup_{0\leq i\leq k-1}d(f(x_{i}),x_{i+1})\leq\delta.$
We say that $f$ is chain transitive if for any $x,y\in X$ and $\delta>0$, there is a $\delta$-chain $(x_{i})_{i=0}^{k}$ of $f$ such that $x_{0}=x$ and $x_{k}=y$.
Remark 1.
For a continuous map $f\colon X\to X$, the following conditions are equivalent
  • $f$ is chain transitive,
  • for any open subset $U$ of $X$, $f(\overline{U})\subset U$ implies $U\in\{\emptyset,X\}$.
Remark 2.
For any continuous map $f\colon X\to X$ and $x\in X$,
$f|_{\omega(x,f)}\colon\omega(x,f)\to\omega(x,f)$
is chain transitive.
Definition 2.
A homeomorphism $H\colon Z\to Z$ of a compact metric space $Z$ is said to be an abstract $\omega$-limit set if there are a homeomorphism $f\colon X\to X$ and $x\in X$ such that
$f|_{\omega(x,f)}\colon\omega(x,f)\to\omega(x,f)$
is topologically conjugate to $H\colon Z\to Z$, i.e., there is a homeomorphism $\phi\colon\omega(x,f)\to Z$ such that $\phi\circ f|_{\omega(x,f)}=H\circ\phi$.
By Remark 1 and [2, Theorem 1], we obtain the following lemma.
Lemma 2.
A homeomorphism $H\colon Z\to Z$ of a compact metric space $Z$ is an abstract $\omega$-limit set if and only if $H$ is chain transitive.
Given a continuous map $f\colon X\to X$, we denote by $M(f)$ the (disjoint) union of minimal sets for $f$. A non-empty closed subset $M$ of $X$ with $f(M)\subset M$ is a minimal set for $f$ if and only if for any closed subset $S$ of $M$, $f(S)\subset S$ implies $S\in\{\emptyset,M\}$. By Zorn’s lemma, there is at least one minimal set for $f$; therefore, $M(f)\neq\emptyset$.
A continuous map $f\colon X\to X$ is said to be proximal if
$\liminf_{i\to\infty}d(f^{i}(x),f^{i}(y))=0$
for all $x,y\in X$. We know that for a continuous map $f\colon X\to X$, the following conditions are equivalent
  • $f$ is proximal,
  • $M(f)$ is a singleton, i.e., there is $p\in X$ such that $f(p)=p$ and $M(f)=\{p\}$.
Remark 3.
If a continuous map $f\colon X\to X$ is proximal, then so is $f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$.
Here we recall the definition of $\omega^{\ast}$-chaos from [6].
Definition 3.
Let $f\colon X\to X$ be a continuous map. A subset $S$ of $X$ is said to be an $\omega^{\ast}$-scrambled set for $f$ if for any $x,y\in S$ with $x\neq y$,
  • $\omega(x,f)\setminus\omega(y,f)$ contains an infinite minimal set for $f$,
  • $\omega(x,f)\cap\omega(y,f)\neq\emptyset$.
We say that $f$ is $\omega^{\ast}$-chaotic if there is an uncountable $\omega^{\ast}$-scrambled set for $f$.
Remark 4.
Given a continuous map $f\colon X\to X$, it is obvious that an $\omega^{\ast}$-scrambled set for $f$ is an $\omega$-scrambled set for $f$, thus if $f$ is $\omega^{\ast}$-chaotic, then $f$ is $\omega$-chaotic. If $f$ is proximal, then every $\omega^{\ast}$-scrambled set $S$ for $f$ satisfies $|S|\leq 1$; therefore, $f$ is not $\omega^{\ast}$-chaotic.
Remark 5.
In Section 6 of [5], an example is given of a continuous map $f\colon X\to X$ such that
  • $X=\overline{Per(f)}$,
  • $f$ is mixing,
  • $f$ is $\omega$-chaotic but not $\omega^{\ast}$-chaotic.
We can now present the promised examples.
Example 1.
Let $Z=S^{1}=\{(a,b)\in\mathbb{R}^{2}\colon a^{2}+b^{2}=1\}$ and let $q=(0,1)$. We consider the standard bijection $i\colon\mathbb{R}\to Z\setminus\{q\}$ defined as
$i(r)=\left(\frac{2r}{r^{2}+1},\frac{r^{2}-1}{r^{2}+1}\right)$
for all $r\in\mathbb{R}$, and define a homeomorphism $H\colon Z\to Z$ by
  • $H(q)=q$,
  • $H(i(r))=i(r+1)$ for all $r\in\mathbb{R}$.
Since $H$ is chain transitive, by Lemma 2, there are a homeomorphism $f\colon X\to X$ and $x\in X$ such that
$f|_{\omega(x,f)}\colon\omega(x,f)\to\omega(x,f)$
is topologically conjugate to $H\colon Z\to Z$, i.e., there is a homeomorphism $\phi\colon\omega(x,f)\to Z$ such that $\phi\circ f|_{\omega(x,f)}=H\circ\phi$. Consider $Y$, $h$, $g$ as in Corollary 1. Letting $p=\phi^{-1}(q)$, we see that
  • $f(p)=p$,
  • $\omega(x,f)$ is an uncountable set,
  • $\omega(x,f)\cap Per(f)=\{p\}$; therefore, $\omega(x,f)\cap Per(f)\neq\emptyset$ and $\omega(x,f)\setminus Per(f)\neq\emptyset$.
It follows that $g$ is $\omega$-chaotic. Since $M(h)=\{p\}$, a singleton, $h$ is proximal and so is $g$. It follows that $g$ is not $\omega^{\ast}$-chaotic. It is obvious that the topological entropy of $g$ is zero.
A continuous map $f\colon X\to X$ is said to be
  • weakly mixing if $f\times f\colon X\times X\to X\times X$ is transitive,
  • mixing if for any non-empty open subsets $U,V$ of $X$, there is $j\geq 0$ such that $f^{i}(U)\cap V\neq\emptyset$ for all $i\geq j$.
We say that $f$ is uniformly rigid if for any $\epsilon>0$, there is $j\geq 1$ such that
$\sup_{x\in X}d(x,f^{j}(x))\leq\epsilon.$
Let $g=f^{\mathbb{N}}\colon X^{\mathbb{N}}\to X^{\mathbb{N}}$ and note that
  • as a consequence of Furstenberg’s $2$ implies $n$ theorem (see [4, Proposition II.3]), if $f$ is weakly mixing, then so is $g$,
  • if $f$ is mixing, then so is $g$,
  • if $f$ is uniformly rigid, then so is $g$.
Example 2.
Given a transitive continuous map $f\colon X\to X$ such that $X$ is an infinite set, we fix $x\in X$ with $X=\omega(x,f)$. Let $Y$, $h$, $g$ be as in Corollary 1. Note that, in this case, $Y=X$ and $h=f$.
  • (1)
    If $Per(f)\neq\emptyset$, in particular, if $f$ is Devaney chaotic, then $g$ is $\omega$-chaotic.
  • (2)
    As an example given in Section 5 of [3], if $f$ is proximal, weakly mixing, and uniformly rigid, then $g$ is proximal, weakly mixing, uniformly rigid, $\omega$-chaotic, and not $\omega^{\ast}$-chaotic.
  • (3)
    As an example constructed in Section 4 of [10], if $f$ is proximal and mixing, then $g$ is proximal, mixing, $\omega$-chaotic, and not $\omega^{\ast}$-chaotic.
Example 3.
Let $Z=S^{1}=\{(a,b)\in\mathbb{R}^{2}\colon a^{2}+b^{2}=1\}$. We define a homeomorphism $H\colon Z\to Z$ by $H(c)=c$ for all $c\in Z$. Since $H$ is chain transitive, by Lemma 2, there are a homeomorphism $f\colon X\to X$ and $x\in X$ such that
$f|_{\omega(x,f)}\colon\omega(x,f)\to\omega(x,f)$
is topologically conjugate to $H\colon Z\to Z$. Note that $f(p)=p$ for all $p\in\omega(x,f)$. Consider $Y$, $h$, $g$ as in Corollary 1. As in the proof of Theorem 1, letting $p\in\omega(x,f)$ and $\Lambda=\{p\}$, we have a Cantor set $S$ in $Y^{\mathbb{N}}$ such that for any $s,t\in S$ with $s\neq t$,
  • $\omega(s,g)\setminus\omega(t,g)$ is an uncountable set,
  • $(p,p,p,\dots)\in\omega(s,g)\cap\omega(t,g)$ and so $\omega(s,g)\cap\omega(t,g)\neq\emptyset$.
However, since $g(v)=v$ for all $u\in Y^{\mathbb{N}}$ and $v\in\omega(u,g)$, $g$ is not $\omega$-chaotic.